OpenStudy (anonymous):

solve the equations to find r,s,and t. 2r+2+t=14 -r-3s+2t=-2 4r-6s+3t=-5

OpenStudy (anonymous):

by which method?

OpenStudy (anonymous):

i prefer elimination, though ill take which ever is the easiest since elimination hasnt been working for me.

OpenStudy (anonymous):

ok, i have to go and pray right now. i will be back in 20 mins and i will walk you through it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it 2r+2s+t=14 or 2r+2+t=14

OpenStudy (anonymous):

sorry it is 2r+s+t=14

OpenStudy (anonymous):

i dont know what im doing ive tried the problem several times and my answers are always different and not the answer it is supposed to be (1,5,7) though i had another problem too and i got it.

OpenStudy (anonymous):

are you getting it ok?

OpenStudy (anonymous):

im working on it on a word doc

OpenStudy (anonymous):

oh ok! that will be handy!

OpenStudy (anonymous):

i have to go but i will still have the question open so thankyou very much in advance for helping me.

OpenStudy (anonymous):

thank You! @surdawi

OpenStudy (anonymous):

dont forget my medal lol